# thanks: equal contributionthanks: equal contribution

### v1

@@ Abstract
We investigated the optimal control of a continuous variable system, focusing on entanglement generation in an optomechanical system without utilizing Fock basis cutoffs. Using the Krotov algorithm to optimize the dynamics of the covariance matrix, we illustrated how to design a control objective function to manipulate the dynamics of the system to generate a desirable target state. We showed that entanglement between the macroscopic mechanical mirror and the quantum optical cavity can be reliably generated through imposing the control on the detuning of the external laser field. It has be shown that the control may be still achieved when imposing spectral constraints on the external field to restrict it to low-frequency components. In addition, we systematically studies the effects of quantum control on non-Markovian open system dynamics. We observed that memory effects can play a beneficial role in mitigating the detrimental impact of environmental noises. Specifically, the entanglement generated shows reduced decay in the presence of these memory effects.
@@ I Introduction
To date, there is a wide range of quantum control protocols available [40, 41, 42, 43, 44], such as stimulated Raman adiabatic passage (STIRAP) [45], chopped random basis [46, 47] and gradient ascent methods [48]. Optimal controls using machine learning tools have also been recently proposed [49, 50, 51]. Among the various control strategies, the Krotov control [52, 53, 54, 55, 56] is a versatile algorithm that can be applied to Bose-Einstein condensate [57], photochemistry [54, 58], non-linear Schrödinger equation [57, 53], and optimal control under open systems [59], as well as controlling towards entanglers [60, 61]. For optomechanical system, controls have been proposed using continuous measurement [62] and feedback control [63]. Entanglement generation using reservoir engineering [64], pulsed interaction [65], via dissipative effects [66] and external pulses [67] have also been reported. However, to implement a numerical implementation for the optimal control of continuous variable (CV) systems [35, 68, 69] would generally require some sort of cut-off, such as cut off the Fock state basis larger than some photon number [70, 71, 72, 73]. Such cutoffs may either difficult to estimate, or introduces errors due to the cutoff. Based on the initial state and target state of the control, this may also restrict the states available for study where only states with some low photon numbers or in a small subspace is available. It can also pose issues in many-body systems as the number of particles grows due to the corresponding exponential growth of the Hilbert space.
@@
In this paper we show how to use the Krotov method to control CV systems, specifically focusing on entanglement generation in optomechanical systems. To address the limitations associated with Fock-basis truncation, we propose applying the Krotov control method to the expectation value equations. Specifically, we adapt the Krotov control objective function for the equation of motion of the quantum system’s covariance matrix, which directly affects the degree of entanglement. This approach allows for the study of quantum system control without the constraints imposed by Fock-basis truncation. We also investigate key properties of the control algorithm, such as convergence behavior and spectral limits, to ensure its efficacy in practical applications. A realistic description of quantum systems necessitates a more general treatment of open systems, where the environmental memory must be considered. We derive the equation of motion governing the covariance matrix in a general non-Markovian open system setting, and examine the robustness of optimal control strategies under these dynamics. The general approach outlined in this paper will be useful for the study of controlling various continuous variable (CV) quantum systems beyond the specific examples discussed. The paper is organized as follows: We first introduce a general outline of how to utilize the Krotov control algorithm on the CV system in Sec. II, then turn to a more specific example of generating entanglement between the mechanical mode and the quantum cavity in an optomechanical system in Sec. III, with a study on the non-Markovian open systems in Sec. IV.
@@ II Optimal control of CV systems
can drive the system dynamics to some target states at a prescribed time $t_{f}$, where $H_{0}$ and $H_{i}$ are time-independent Hamiltonian describing the uncontrolled part and components under control, respectively. To facilitate control of the dynamics of quantum systems, many different protocols have been proposed [40, 41, 42, 43, 44, 45, 46, 47]. A common problem facing quantum optimal controls is that there would be an interdependence of the control field and the quantum state. Without loss of generality, we consider here the case of a single control target. Among the various control strategies, one outstanding algorithm is the Krotov method [52, 53, 54, 55, 56]. By clever mathematical construction, the Krotov control method decouples the interdependence of the control field and the quantum state under control, and uses an iterative algorithm to minimize a target function of the form
@@
In order to drive an initial states to a target state, the target function $J_{T}$ is usually taken as the infidelity $1-f$ with $f$ being the fidelity between the target state and the evolved state under control. However, with continuous variable systems, one issue is that to facilitate the numerical calculations, the Hamiltonian and the quantum states needs to have some cutoff in the Hilbert space. While in certain cases the cutoff $N_{c}$ in the Fock space can be estimated, it can still pose issues where such estimation is difficult. One complication lies in that while the cutoff may be estimated for the initial and final states, special care needs to be taken such that this cutoff applies during all times in the evolution, which can be non-trivial to impose. Moreover, especially with many-body systems, if the required cutoff is large, the computation cost can get very expensive, as the Hilbert space grow exponentially with the number of particles $n$ considered as $N_{c}^{n}$. This is especially relevant for tasks such as entanglement generation in CV systems, where operators not conserving the total number of excitations such as non-linear interactions are called for [79, 36, 80, 81].
@@ III Entanglement generation in an optomechanical system
where $a$ and $b$ are the quantum fluctuations of optical and mechanical modes around their mean values ($\alpha$, $\beta$), with the linearization $\tilde{a}=a+\alpha$ and $\tilde{b}=b+\beta$. $G=\alpha g$ is the effective coupling rate, and $\Delta=\omega-\omega_{c}+2G^{2}/\omega_{m}$ is the modified detuning. To generate entanglement between the mechanical mirror and the optical cavity, we propose to control the system by tuning the frequency of the external laser, which may be realized experimentally [84, 85]. This would lead to a time-dependent detuning parameter which would serve as the control parameter $f(t)$, with the Hamiltonian
@@
$=\frac{f(t)}{2}\left[p_{c}^{2}+q_{c}^{2}\right]+\frac{\omega_{m}}{2}\left[p_{m}^{2}+q_{m}^{2}\right]+2Gq_{c}q_{m}$ (12)
@@
and $\mathcal{M}_{c}=\mathrm{diag}(1,0,1,0)$. The target CM is chosen to be the entangled two-mode squeezed state [86, 69] given by $\exp\left[r\left(ab-a^{\dagger}b^{\dagger}\right)\right]|00\rangle$ with a squeezing parameter $r\in R$, where $|00\rangle$ is the two-mode vacuum state. The explicit expression of the two-mode squeezed state CM can be written in the form of
@@
Choosing system parameters $\omega_{m}=1$, $G=0.1$, we now use the Krotov control algorithm to generate the control field to drive an initial vacuum state $|00\rangle$ to the target desired two-mode squeeze state with a squeezing parameter $r=1.25$. The target negativity $N_{T}\approx 3.6067$, and the control goal is set to be such that the distance between the evolved CM and the target CM is less than $10^{-4}$, with a control run-time of $t_{f}=60$. The initial guess for the control is set to be a constant $f(t)=0$. The resulting controlled detuning $f(t)$ is shown in Fig. 1 (a) as the solid blue line. We also show the entanglement dynamics between the mechanical mirror and the optical cavity in Fig. 1 (c), as the ratio of the negativity of the evolved state divided by the target negativity (solid blue line). It can be seen that the controlled detuning is on the same order-of-magnitude with other parameters of the Hamiltonian, and we are able to reach the target entangled state at the end of the evolution.
@@
III.1 Krotov control with optional spectral cutoff
In practical scenarios, it is often preferable to use a control field that is neither too strong in magnitude nor oscillates too rapidly. While it can be observed clearly in Fig. 1 (a), we may also quantitatively study it by taking a Fourier transformation of the control field. Given that the control field is real-valued, we use the standard discrete cosine transformation (DCT),
@@ III.1 Krotov control with optional spectral cutoff
IV Open system effects on the optimal controls
@@
In this section, we will study how quantum entanglement is affected by the environment effects. It is known that entanglement is a fragile resource and may be susceptible to many environment effects such as the decoherence, dephasing and entanglement sudden death [92, 93]. To study how the proposed control strategy is robust againt the influences of open system effects, we consider the optomechanical system embedded in a bosonic environment,
@@
where $L$ describes the damping of the mechanical oscillator and $g_{j}$ is the coupling strength between the system and the environment. To obtain the reduced system dynamics, we use the quantum state diffusion (QSD) equation [94, 95, 26, 96], which projects the bath modes onto the Bargmann coherent states to arrive at a stochastic Schrödinger equation for a pure representing a quantum trajectory,
@@ V Conclusion
As an illustrative example, we applied the control strategy to generate quantum entanglement in an optomechanical system between a mechanical mirror and an optical cavity. Our control protocol reliably drive an initial vacuum state into an entangled two-mode squeezed state. Furthermore, the protocol exhibits good convergence behavior, with the distance between the controlled state and the target state converging exponentially with the number of Krotov iterations. The resulting control field is stable, exhibiting neither excessive amplitude nor high-frequency oscillations. We also imposed spectral constraints and found that satisfactory control fields can still be achieved within these limits. Finally, we have studied the effects of a general non-Markovian environment on the generated entanglement. The equation of motion for the CM under non-Markovian dynamics shows that quantum memory effects can preserve the entanglement, resulting in slower decay compared to a memoryless (Markovian) bath. The protocol and methodology presented here are general and adaptable to other continuous variable systems, offering insights into optimal control strategies and the impact of open system dynamics.

### v2

@@ Abstract
We investigate the optimal control of a continuous variable (CV) system, focusing on entanglement generation in an optomechanical system without utilizing Fock basis cutoffs. Using the Krotov algorithm to optimize the dynamics of the covariance matrix, we illustrate how to design a control objective function to manipulate the dynamics of the system to generate a desirable target state. We showed that entanglement between the macroscopic mechanical mirror and the quantum optical cavity can be reliably generated through imposing the control on the detuning of the external laser field. It has been shown that the control may still be achieved when imposing spectral constraints on the external field to restrict it to low-frequency components. In addition, we systematically study the effects of quantum control on non-Markovian open system dynamics. We observed that memory effects can play a beneficial role in mitigating detrimental impacts of environmental noises. Specifically, the generated entanglement shows a reduced decay rate in the presence of these memory effects.
@@ I Introduction
To date, there is a wide range of quantum control protocols available [40, 41, 42, 43, 44], such as stimulated Raman adiabatic passage (STIRAP) [45], chopped random basis [46, 47] and gradient ascent methods [48]. Optimal controls using machine learning tools have also been recently proposed [49, 50, 51]. Among the various control strategies, the Krotov control [52, 53, 54, 55, 56] is a versatile algorithm that can be applied to Bose-Einstein condensate [57], photochemistry [54, 58], non-linear Schrödinger equation [57, 53], and optimal control under open systems [59], as well as controlling towards entanglers [60, 61]. For optomechanical system, controls have been proposed using continuous measurement [62] and feedback control [63]. Entanglement generation using reservoir engineering [64], pulsed interaction [65], via dissipative effects [66] and external pulses [67] have also been reported. However, to implement a numerical implementation for the optimal control of continuous variable (CV) systems [35, 68, 69] would generally require some sort of cut-off, such as cut off the Fock state basis larger than some photon number [70, 71, 72, 73]. Such cutoffs are difficult to estimate and may introduce errors due to the imposed limitations. Depending on the initial state and the target state of the control, this approach can also restrict the range of accessible states, confining the study to states with low photon numbers or those residing within a small subspace. In addition, it can also pose issues in many-body systems as the number of particles grows due to the corresponding exponential growth of the Hilbert space.
In this paper we show how to use the Krotov method to control and manipulate a CV system, in focusing on entanglement generation in optomechanical systems. Specifically, to address the limitations associated with Fock-basis truncation, we propose applying the Krotov control method to the equations of expectation values. We adapt the Krotov control objective function for the equation of motion of the quantum system’s covariance matrix, which directly affects the degree of entanglement. This approach allows for the study of quantum system control without the constraints imposed by Fock-basis truncation.
@@
In this paper, we demonstrate the application of the Krotov method to control and manipulate a CV system, focusing specifically on entanglement generation in optomechanical systems. To address the limitations associated with Fock-basis truncation, we propose applying the Krotov control method to the equations of expectation values. By adapting the Krotov control objective function for the equation of motion of the quantum system’s covariance matrix, which directly impacts the degree of entanglement, our approach allows for the study of quantum system control without the constraints imposed by Fock-basis truncation. Moreover, we also investigate key properties of the control algorithm, such as convergence behavior and spectral limits, to ensure its efficacy in practical applications. A realistic description of quantum systems necessitates a more general treatment of open systems, where the environmental memory must be considered. We derive the equation of motion governing the covariance matrix in a general non-Markovian open system setting, and examine the robustness of optimal control strategies under these dynamics. The general approach outlined in this paper will be useful for the study of controlling various continuous variable (CV) quantum systems beyond the specific examples discussed. The paper is organized as follows: We first introduce a general outline of how to utilize the Krotov control algorithm on the CV system in Sec. II, then turn to a more specific example of generating entanglement between the mechanical mode and the quantum cavity in an optomechanical system in Sec. III, with a study on the non-Markovian open systems in Sec. IV.
@@ II Optimal control of CV systems
can drive the system dynamics to some target states at a prescribed time $t_{f}$, where $H_{0}$ and $H_{i}$ are time-independent Hamiltonian describing the uncontrolled part and components under control, respectively. To facilitate control of the dynamics of quantum systems, many different protocols have been proposed [40, 41, 42, 43, 44, 45, 46, 47]. A common problem facing quantum optimal controls is that there would be an interdependence of the control field and the quantum state. Without loss of generality, we consider here the case of a single control target. Among the various control strategies, one outstanding algorithm is the Krotov method [52, 53, 54, 55, 56]. In the context of continuous variable systems, this technique has recently been utilized to generate single-mode squeezing [78] and engineer beam splitter interactions [79]. By clever mathematical construction, the Krotov control method decouples the interdependence of the control field and the quantum state under control, and uses an iterative algorithm to minimize a target function of the form
@@
In order to drive an initial states to a target state, the target function $J_{T}$ is usually taken as the infidelity $1-F$ with $F$ being the fidelity between the target state and the evolved state under control. However, with continuous variable systems, one issue is that to facilitate the numerical calculations, the Hamiltonian and the quantum states needs to have some cutoff in the Hilbert space. While in certain cases the cutoff $N_{c}$ in the Fock space can be estimated, it can still pose issues where such estimation is difficult. One complication lies in that while the cutoff may be estimated for the initial and final states, special care needs to be taken such that this cutoff applies during all times in the evolution, which can be non-trivial to impose. Moreover, especially with many-body systems, if the required cutoff is large, the computation cost can get very expensive, as the Hilbert space grows exponentially with the number of particles $n$ considered as $N_{c}^{n}$. This is especially relevant for tasks such as entanglement generation in CV systems, where operators not conserving the total number of excitations such as non-linear interactions are called for [81, 36, 82, 83].
@@ III Entanglement generation in an optomechanical system
Figure 1: Schematic of the model under consideration. The single-sided optomechanical system consists of a mechanical mirror mode coupled to an optical cavity mode driven by an external coherent laser. The laser is tunable and set to dynamically change the effective detuning $\Delta$ to realize the optimal control.
@@
where $a$ and $b$ are the quantum fluctuations of optical and mechanical modes around their mean values ($\alpha$, $\beta$), with the linearization $\tilde{a}=a+\alpha$ and $\tilde{b}=b+\beta$. $G=\alpha g$ is the effective coupling rate, and $\Delta=\omega-\omega_{c}+2G^{2}/\omega_{m}$ is the modified detuning. The fluctuations can be determined by [13, 12, 26] as $\left[i(\omega-\omega_{c})-ig(\beta+\beta^{*})-\kappa_{a}\right]\alpha=i\Omega_{d}$ and $-i\omega_{m}\beta=ig|\alpha|^{2}$ with classical cavity leakage rate as $\kappa_{a}$. This represents possibly one of the simplest models for the optomechanical cavity. To generate entanglement between the mechanical mirror and the optical cavity using the optimal control strategy outlined in the last section, we propose to use a tunable drive laser on the optomechanical system, which may be realized experimentally [86, 87]. We show a schematic of the controlled model under consideration in Fig. 1. Specifically, by tuning the frequency and amplitude of the external drive laser, one may dynamically change the detuning $\Delta$ in Eq. (11). Denoting the now time-dependent detuning parameter which would serve as the control function $f(t)$, we have
@@
$\simeq\frac{f(t)}{2}\left[p_{c}^{2}+q_{c}^{2}\right]+\frac{\omega_{m}}{2}\left[p_{m}^{2}+q_{m}^{2}\right]+2Gq_{c}q_{m},$ (12)
@@
and $\mathcal{M}_{c}=\mathrm{diag}(1,0,1,0)$.
Here, the target CM is chosen to be the entangled two-mode squeezed state [88, 69] given by $S(r)|00\rangle=\exp\left[r\left(ab-a^{\dagger}b^{\dagger}\right)\right]|00\rangle$ with a squeezing parameter $r\in R$, where $|00\rangle$ is the two-mode vacuum state. The explicit expression of the two-mode squeezed state CM can be written in the form of
@@
Choosing system parameters $\omega_{m}=1$, $G=0.1\omega_{m}$, we now use the Krotov control algorithm to generate the control field to drive an initial vacuum state $|00\rangle$ to the target desired two-mode squeezed state with a squeezing parameter $r=1.25$. Experimentally, the single photon coupling $g_{0}$ has been realized between $10^{-6}\omega_{m}$ and $10\omega_{m}$ [13]. Thus, our choice of the effective coupling rate $G=0.1\omega_{m}$ could be made within experimental viable ranges, where $G=g_{0}\sqrt{n_{c}}$ and $n_{c}$ is the cavity photon number. Notably, coupling rates of similar values has been reported in [108, 109, 110] at $G\sim 0.12\omega_{m}$, $G\sim 0.2\omega_{m}$ and $G\sim 0.34\omega_{m}$. The target negativity $N_{T}\approx 3.6067$, and the control goal is set to be such that the distance between the evolved CM and the target CM is less than $10^{-4}$, with a control run-time of $t_{f}=60$. The initial guess for the control is set to be a constant $f(t)=0$. The resulting controlled detuning $f(t)$ is shown in Fig. 2 (a) as the solid blue line. We also show the entanglement dynamics between the mechanical mirror and the optical cavity in Fig. 2 (c), as the ratio of the negativity of the evolved state divided by the target negativity (solid blue line). It can be seen that the controlled detuning is on the same order-of-magnitude with other parameters of the Hamiltonian, and we are able to reach the target entangled state at the end of the evolution.
@@
As a study of the robustness of the control, we consider two distinct cases: for weak coupling rates $G\ll\omega_{m}$, if one lets $f(t)=-\omega_{m}$ we have an effective Hamiltonian, after rotating out $-\omega_{m}a^{\dagger}a+\omega_{m}b^{\dagger}b$, and applying the rotating wave approximation (RWA), $H^{\prime}\approx G(ab+h.c.)$ in the interaction picture, equivalent to the two-mode squeezing operator. Going back to the Schrödinger picture, the entanglement measured by the logarithm negativity would grow linearly without control (upper panel of Fig. 4, orange dashed line, with a constant $f(t)=-\omega_{m}$). Note that in this case, the Krotov control is still valid since we are able to reach the target state at the prescribed runtime $t_{f}=30$ for a target squeezing factor $r=0.2$ (blue solid line, upper panel of Fig. 4). On the other hand, when the coupling strength $G$ is larger, the RWA would no longer apply. In such case, the uncontrolled entanglement dynamics may just oscillate while a much higher degree of entanglement can be reached with the Krotov control applied (squeezing factor $r=1.0$, lower panel of Fig. 4, where for the uncontrolled case we set a constant $f(t)=0$). It’s worth pointing out that for the uncontrolled dynamics in the RWA case, while the entanglement approximately grows linearly in time as the effective squeezing parameter under $H^{\prime}$ is $Gt$, the linear growth would be limited to a very slow rate since we need $G\ll\omega_{m}$ for the RWA to hold in this uncontrolled scenario.
Figure 4: Dynamics in different parameter regions with (blue solid line) and without control (orange dashed line), for a total runtime of $t_{f}=30$: upper panel is for the case where $G/\omega_{m}=0.01\ll 1$ with a target squeezing factor $r=0.2$ and the lower panel is for $G/\omega_{m}=0.1$, with a target squeezing factor $r=1.0$. The green dotted line denotes the target entanglement.
In the control of quantum dynamics, one is often interested in how to achieve the control target in the shortest amount of time, a topic generally known as the quantum speed limits (QSL) [93, 94, 95]. With its roots in the uncertainty principle, QSL has been rediscovered in various studies such as geometric approaches [96]. In the context of optimal controls [97], one may estimate the quantum speed limit time as [98]
$T_{\rm QSL}\sim\Delta E_{0}^{-1}\arccos|\langle\psi(0)|\psi_{T}\rangle|,$ (20)
where $|\psi_{T}\rangle$ is the target state and $|\psi(0)\rangle$ is the initial state. For the system under consideration here, the variance $\Delta E_{0}=\sqrt{\langle H^{2}\rangle-\langle H\rangle^{2}}=G$ for the initial vacuum state and $H=H_{c}$ in Eq. (12) with an initial guess control $f(t)=0$ as with the vacuum initial state the term $a^{\dagger}a$ has no contribution. The target two-mode squeezed state can be written in the Fock basis as [88, 99] $S(r)|00\rangle=(\cosh r)^{-1}\sum_{n=0}^{\infty}(-\tanh r)^{n}|n,n\rangle$ so the inner product between the initial and target state is $1/\cosh r$. The QSL time is thus estimated as $T_{\rm QSL}\sim\arccos(1/\cosh r)/G$. In Fig. 5 we show the relative final entanglement $N(t_{f})/N_{T}$ where $N_{T}$ is the target negativity, as a function of runtime $t_{f}$ and the coupling strength $G$, as well as a boundary estimated by the QSL Eq. (20). Here, the target state is chosen to be a two-mode squeezed state with a squeezing parameter $r=0.8$. It can be seen that one can reach the target entangled state whenever the evolution time is larger than the limit imposed by the QSL, while the target state may not be reachable when the evolution time is shorter than the QSL time. In this sense, with the control protocol presented here, we can have an optimal control at the quantum speed limit.
Figure 5: Final relative entanglement $N(t_{f})/N_{T}$ where $N_{T}$ is the target negativity, as a function of the coupling strength $G$ and runtime $t_{f}$. The gray dashed line denotes analytical estimation of the quantum speed limit time. The target state is chosen to be a two-mode squeezed state with a squeezing parameter $r=0.8$.
Krotov control with optional spectral cutoff
In a typical practical scenario, it is often preferable to use a control field that is neither too strong in magnitude nor oscillates too rapidly. While it can be observed clearly in Fig. 2 (a), we may quantitatively study it by taking a Fourier transformation of the control field. Given that the control field is real-valued, we use the standard discrete cosine transformation (DCT),
@@
IV Open system effects on the entanglement generation
@@
In this section, we will study how the optimal control of quantum entanglement is affected by the environment effects. It is well-known that entanglement is a fragile resource, highly susceptible to various environmental effects such as decoherence, dephasing, and entanglement sudden death [101, 102]. To study how the proposed control strategy is robust against the influences of open system effects, we consider the optomechanical system embedded in a bosonic environment,
@@
where $L$ describes the damping of the mechanical oscillator and $g_{j}$ is the coupling strength between the system and the environment. To obtain the reduced system dynamics, we use the quantum state diffusion (QSD) equation [103, 104, 26, 105], which projects the bath modes onto the Bargmann coherent states to arrive at a stochastic Schrödinger equation for a pure state representing a quantum trajectory,
@@ V Conclusion
As an illustrative example, we applied the control strategy to generate quantum entanglement in an optomechanical system between a mechanical mirror and an optical cavity. Our control protocol reliably drive an initial vacuum state into an entangled two-mode squeezed state. Furthermore, the protocol exhibits good convergence behavior, with the distance between the controlled state and the target state converging exponentially with the number of Krotov iterations. The resulting control field is stable, exhibiting neither excessive amplitude nor high-frequency oscillations. We also imposed spectral constraints and found that satisfactory control fields can still be achieved within these limits. We also study the quantum speed limits and found the control is achieved when the runtime is larger than the limit imposed by the QSL. Finally, we have studied the effects of a general non-Markovian environment on the generated entanglement. The equation of motion for the CM under non-Markovian dynamics shows that quantum memory effects can preserve the entanglement, resulting in slower decay compared to a memoryless (Markovian) bath. Our choice of the effective coupling rate $G=0.1\omega_{m}$ could be feasible with current technologies [13, 108, 109, 110], and the protocol on the control of detuning could be a more experimentally viable choice [86, 87]. Note that theoretically, since detuning can be both positive and negative, it provides greater flexibility in calculating the control fields. In contrast, the laser intensity or coupling must be modulated to remain non-negative, which introduces an additional control constraint. Note in this case it is still possible to realize the desired control when the coupling is physically permissible to the control mechanism. While our protocol allows the system to achieve the target entangled state within a predetermined time, it cannot preserve the target states indefinitely. One way to achieve long-term preservation of the target state is to have the ability to control the coupling $G$ (e.g., turn off or set the coupling $G$ to 0). In this case, the effective Hamiltonian would be $\omega_{m}b^{\dagger}b-\Delta a^{\dagger}a$, which would merely add a phase to the squeezing parameter and keep the negativity (entanglement) constant after $t_{f}$. The protocol and methodology presented here are general and adaptable to other Gaussian continuous variable systems [*[Afterthecompletionofthiswork, webecameawareofthisrecentworkusingsimilartechniquesforchargingofquantumbatteries:][]Rodriguez2024a], offering insights into optimal control strategies and the impact of open system dynamics.